# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Bridges

### v1

@@ Abstract
Between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen via attacks on “bridges” — decentralized services designed to allow inter-blockchain exchange. While the individual exploits in each attack vary, a single design flaw underlies them all: the lack of end-to-end value accounting in cross-chain transactions. In this paper, we empirically analyze twenty million transactions used by key bridges during this period. We show that a simple invariant that balances cross-chain inflows and outflows is compatible with legitimate use, yet precisely identifies every known attack (and several likely attacks) in this data. Further, we show that this approach is not only sufficient for post-hoc audits, but can be implemented in-line in existing bridge designs to provide generic protection against a broad array of bridge vulnerabilities.
1 Introduction
@@
Careful accounting is key to the correct management of assets in virtually all financial systems. Indeed, since Paciloli popularized double-entry bookkeeping in the 14th century, it has been standard practice to separately track inflows (credits) and outflows (debits) to establish a net position — a balance sheet.
@@
Such accounting is implicit in blockchains as each and every transaction is recorded, immutable and in possession of implicit integrity. Any change in ownership for a given token is explicitly recorded via some past transaction and thus the net position for assets in a blockchain is generally consistent and well-known. However, these integral properties only hold within a single blockchain. As soon as one wishes to engage in transactions between blockchains (e.g., trading Ethereum on Solana), it requires building a system that steps outside these isolated environments and implements its own financial calculus between them, including its own accounting.
@@
Today, one of the principal mechanisms for such transactions is the cross-chain bridge — a service using a pair of “smart contracts” (immutable programs stored on a blockchain) to synthesize inter-chain transactions that are not possible to express natively on a single chain. [^1] Some newer blockchain ecosystems, such as Cosmos and Avalanche, are themselves composed of multiple blockchains and offer some support for bridging across their own chains, but not with external chains. Such bridges are an extremely popular component of what is commonly referred to as “decentralized finance” (DeFi), and they routinely manage transaction volumes with value equivalent to \$US8–10 billion per month [7].
@@
However, bridges are just code and they can have bugs in their implementations, in the other services they depend upon, or in the mechanisms used to guard the integrity of their cryptographic secrets. As a result, attackers can, and do, exploit such vulnerabilities to extract significant value from such systems. For example, between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen in an array of attacks on bridges. In this work, we identify the lack of explicit accounting as the shared and signal failure across virtually all of these attacks. Considering the twelve largest attacks (each \$US1m or more) during this period, the details of their vulnerabilities vary considerably, but all share the property that transactions are allowed to be “unbalanced” (i.e., that outflow can exceed inflow, less costs). We show that auditing transactions for this property is sufficient to retrospectively and automatically identify each of these past attacks. Moreover, such audits can be performed in real-time — either to alert bridge operators about potential fraud, or prevent such unbalanced transactions from completing altogether.
In particular, our paper makes three main contributions:
  • $\triangleright$
    Balance Invariant. We introduce the notion of a simple balance invariant, outflow = inflow - costs, which can be calculated automatically from the transaction formats used by current bridges. We hypothesize that this invariant should naturally hold for benign transactions, but not for those exploiting bridge vulnerabilities.
  • $\triangleright$
    Retrospective validation. We validate our hypothesis by empirically auditing over 20 million past transactions (across more than 20 blockchains, between 11 different bridges) and show that the balance invariant is sufficient to identify each and every attack for which we have collected ground truth. Moreover, we identify vanishingly few other transactions that violate this invariant, virtually all of which deserve further scrutiny. Some are clear contract implementation errors, others are likely new, unreported thefts, and yet others capture bridge behaviors that are deserving of considerably more transparency (e.g., unilateral issuance of millions of dollars worth of uncollateralized tokens).
  • $\triangleright$
    Real-time auditing. We show our approach is not only useful for retrospective analysis, but can be used to audit bridge transactions and detect new attacks in real-time. We implement and demonstrate an online audit of the Wormhole bridge and show that it automatically detects every attack we inject. Further, we introduce and implement a proof-of-concept for a new bridge architecture, called announce-then-execute, that incorporates such auditing into the transaction flow itself — thereby preventing unbalanced malicious bridge transactions from ever completing. Our approach treats the most complicated components of bridges (e.g., verifying relayed messages) as black boxes, adds no new attack surface for theft, and requires minimal changes to existing codebases — a benefit which which we demonstrate by implementing a modified version of the Wormhole bridge.
@@
We argue that this approach is powerful both due to its simplicity (in a legitimate financial transaction, the value paid should be equivalent to the value received, less costs) and its independence from the vagaries of smart contract details. Moreover, it represents a straightforward mechanism that can prevent an entire class of existing attacks on crypto bridges.
@@
Figure 1: Cross-chain token bridging and the different steps attackers can exploit to withdraw unbacked deposits.
2 Background
@@
In this section we first give a brief introduction to smart contract blockchains and cross-chain bridges. We then describe how bridges, under attack, collapse in practice.
2.1 Smart-contract Blockchains
@@
Modern DeFi protocols (e.g., decentralized exchanges and lending protocols) are built on top of “smart-contract” blockchains, like Ethereum. At their core, these chains, like the original Bitcoin blockchain, are distributed ledgers that manage accounts (public keys) and their balances (in native tokens like Ethereum’s ETH). Users interact with these chains by signing and broadcasting (to the distributed nodes that make up the blockchain network) transactions that, for example, transfer funds from their account (using the corresponding account private key) to another user’s account.
@@
Ethereum, and the smart-contract chains it inspired since its release in 2015, differ from Bitcoin by extending the “simple” distributed ledger with a smart contract execution layer. A smart contract on Ethereum is an internal account—and, like a normal, externally owned account (EOA), it has a balance—that has associated code (EVM bytecode), which implements the smart contract’s program logic, and storage, which persists the program’s state across executions. Users interact with (i.e., execute) smart contracts much like they do when transferring funds from their account: they sign a transaction that encodes the smart contract to call (i.e., the contract address), the particular function to execute, and the arguments to call the function with. Instead of simply transferring funds from the user’s account, then, executing such a smart-contract call transaction amounts to executing the smart contract bytecode—and any smart contracts the contract itself calls.
@@
Figure 2: Simplified USDC ERC-20 Token Contract.
2.2 ERC-20 Tokens
@@
One of the immediate applications of smart contracts—and to date still one of most popular—is to create custom tokens (or coins). To launch a new token, an organization no longer needs to launch a new blockchain; they can instead deploy a new contract that implements the ERC-20 token standard interface on a chain like Ethereum and take advantage of existing on-chain infrastructure like decentralized exchanges that make it easy to, for example, trade one kind of token for another—both native tokens (e.g., ETH) and other tokens.222 Most EVM chains, i.e., chains that use Ethereum’s execution layer, follow Ethereum standards like ERC-20. Most non-EVM chains like Solana (which have different execution models) have similar standards (e.g., SPL tokens in Solana’s case).
@@
Figure 2 shows a simplified variant of one such token contract—the USDC stablecoin contract. This contract tracks how many USDC tokens an account has and governs the spending of these tokens (much like a bank governs bank notes). For example, the contract’s mint function lets Circle (the company that owns the USDC contract) mint new tokens into a user’s account—e.g., after receiving the corresponding payment from the user off-chain (in US dollars, as USDC tokens are pegged to the US dollar). The contract exposes the ERC-20 interface that lets users (and smart contracts) transfer tokens from their account by simply calling functions like safeTransferFrom, and, in turn, use the tokens in any DeFi protocol (e.g., lending the USDC to different markets, exchanging USDC for ETH, buying NFTs using the USDC, etc.). Finally, the contract’s burn function “burns” a token out of circulation—and emits an event (Figure 3) that Circle’s off-chain code looks for before allowing a user to withdraw the corresponding USD fiat off-chain.
Figure 3: Event Emitted by an ERC-20 Token (USDC).
2.3 Cross-Chain Token Bridges
@@
While smart contracts make it easy to launch new tokens without spinning up new chains, there is no real shortage of new blockchains being deployed (almost weekly). Indeed, the modern blockchain ecosystem is a many-chain ecosystem. Blockchains like Avalanche, Base, and Solana have different design points—from cheaper “gas” execution costs, to higher throughput, lower latency, and different permission models—that make them better suited for different classes of application. And, this has resulted in applications that span many chains and cross-chain infrastructure that ultimately (try to) allow users to, for example, buy Dogwifhat NFTs on Solana using USDC tokens on Ethereum.
@@
Core to these applications and infrastructure—and the focus of our work—is the cross-chain token bridge. At a high level, a cross-chain token bridge makes it possible for a user to “transfer” their tokens from one chain (e.g., their USDC on Ethereum) to another chain (e.g., Solana) and then use the transferred tokens on the destination chain (e.g., on a Solana exchange trading USDC and Dogwifhat). [^3] While some cross-chain bridge do let users transfer one kind of token (e.g., USDC on ETherem) for a completely different kind of token on the destination chain (e.g., Dogwif on Solana), these bridges essentially fuse the cross-chain token bridge with an exchange (or swap). We focus on token bridges not only because they are fundamental to other kinds of bridges, but also because they have higher volume, more liquidity, and more attacks. Since smart contracts cannot make network requests or otherwise access sate outside their own storage, a cross-chain token bridge consists of smart contracts on both the source and destination chains, and off-chain infrastructure that serves to relay the “transfer” call across the two contracts.
@@
As Figure 1 shows, a typical cross-chain token transfer, i.e., a bridge transaction, consists three phases:
1. Deposit (on source chain)
@@
To initiate a cross-chain token transfer, the user first calls the bridge’s contract on the source chain:
@@
// ... validate the ERC-20 token contract
@@
This contract function—a simplified version of the Qubit bridge deposit function—processes the user’s deposit by validating the transfer request, e.g., against a list supported tokens, and then transferring the user’s ERC-20 tokens to the bridge contract—recall contracts are accounts with balances. Then, the function emits an event recording the user’s deposit details (including the deposit ID, the ERC-20 token contract address, the recipient on the destination chain, and value).
2. Off-chain relay
The emitted event is observed by an off-chain relaying component (step 4), which constantly monitors the source blockchain. The relaying component first verifies the authenticity of the deposit event. If the event is authentic, the relaying component then produces a signed receipt endorsing the deposit (step 5) and, typically, stores the receipt off-chain (step 6). Finally, this signed receipt is sent to the bridge’s withdrawal contract on the destination blockchain (step 7). Who submits the receipt varies across bridges—some submit the receipt on the user’s behalf (in these cases, the signed receipt is simply a signed contract-call transaction), while others give users (and anyone willing to pay gas) the signed receipt and they, in turn, submit the receipt to the withdrawal contract to complete the transfer.
3. Withdraw (on destination chain)
The withdrawal contract on the destination chain, first processes the withdraw request by verifying the receipt and transferring the tokens to the recipient (step 8). In (the simplified) Chainswap’s withdraw case, for example, users call:
@@ 3. Withdraw (on destination chain)
This contract function first charges the caller a fee, then verifies the deposit details against receipt—both that the deposit was not already withdrawn and that the receipt signatures are valid—and finally transfers the tokens to the intended recipient. We consider a cross-chain transaction complete when the asset is released to the recipient (step 9) as such.
We expect every cross-chain bridge transaction to uphold the balance invariant: the value (and kind) of the tokens withdrawn—the outflow—should equal the value (and kind) of the tokens deposited—the inflow—minus the charged fees. In practice, they don’t.
2.4 How Bridges Collapse
In practice, attackers exploited bugs in all three components—the deposit contract, the relayer, and the withdraw contract—and stole signing keys to siphon hundreds of thousands of dollars. Figure 1 highlights the precise steps in the cross-chain token transfer that attackers have historically exploited, including:
@@ 2.4 How Bridges Collapse
    Bugs in the deposit contract. In step 2, the bridge contract verifies that it has received the correct amount of assets before emitting an event. Bugs in this verification logic could allow an attacker to deposit a smaller amount of assets than what is recorded by the bridge in the event (step 3). For example, the Qubit’s deposit(..) function (see above) did not properly validate the token address; this allowed an attacker to pass 0 for the token address, which meant the contract function did not actually transfer any funds from the attacker’s account but still emitted a Deposit event which allowed the attacker to withdraw actual tokens on the destination chain [25].
@@
    Bugs in the off-chain deposit verification. In step 5, the relaying component verifies the authenticity of the deposit event emitted by the bridge contract, including whether the event is emitted by the bridge’s designated contract. Bridges that do not correctly verify deposits would allow attackers to withdraw assets that are never deposited.
@@
    Stolen relayer (or submitter) keys. In step 5, the relaying component signs the deposit receipts which are then submitted to the withdrawal contract as evidence of a valid deposit. If the relayer key is compromised (e.g., was the case with the Ronin bridge [28]) the attacker can forge a valid receipt and then withdraw assets that were never deposited by calling withdraw with the forged receipt.
    The same is true for bridges that submit receipts on behalf of users—and essentially restrict the withdraw callers to privileged submitter accounts. The bridge submitter keys (step 8) have similarly been compromised (e.g., as was the case with AnySwap [4]) and used to withdraw unbacked deposits.
@@
    Bugs in the withdraw verification. In step 8, the bridge contract verifies that the messages are signed by the relaying component and have not been replayed. Bugs in this verification logic (e.g., as we saw with Wormhole [3]) have allowed attackers to supply “valid” payloads that were not signed by the relaying component and replay withdrawal requests with valid deposits receipts that have already been withdrawn.
In this paper, we assume an attacker that exploits any of the aforementioned components or otherwise controls the relayer (or submitter) keys. In the next section, we show that this attacker model and our simple balance invariant checking captures the largest attacks on cross-chain bridges that have happened in the past. In Sections 5 we show that monitoring withdrawals and deposits on the source and destination chains can be used to detect similar attacks in the future. Finally, in Section 6 we show a describe a announce-then-execute bridge design that enforces this invariant to prevent attacks before they happen.
We note that while our threat model captures a wide variety of vulnerabilities, it is not exhaustive. As with any detection system, an attack that violates one of our assumptions (e.g., avoids violating the balance invariant by transferring funds off-bridge or not having a withdrawal, subverts the transaction data used to validate the invariant, etc.) might succeed. We similarly consider other smart contract bugs (beyond bugs in deposit and withdrawal functions) and account key compromises out of scope—and instead focus on the cross-chain bridging aspects which are relatively less well understood. As we show later, our model captures the largest attacks on cross-chain bridges that have happened in the past and systems can use it to prevent similar attacks in the future.
3 Approach
The core hypothesis of our work is that value should be conserved within cross-chain transactions. That is, that the value of the asset inflow in such a transaction (i.e., the deposit) should equal the value of the asset outflow (i.e., the withdrawal). In token transfer bridges, this invariant corresponds to a balancing of the inflow tokens and the outflow tokens (less any fees or transaction costs incurred by the bridge itself). When this balance invariant does not hold it allows a range of opportunities for fraud, all of which involve greater outflows (withdrawals) than inflow (deposits). Figure 4 illustrates such an outcome by graphing the aggregate difference between bridge inflow (via Ethereum) and outflow (via Solana) leading up to and during the February 2022 attack on the Wormhole bridge. The net difference is consistently near zero, with only short positive deviations (representing delayed withdrawals) until the attack in January, at which point there are significant withdrawals without matching deposits—producing a large negative difference.
Testing this invariant on a per transaction basis, is straightforward in principle, but since bridge transaction formats are not standardized, it requires a range of per-bridge and per-chain parsing in practice. Our methodology for normalizing this information focuses on two key pieces of information: a) identifying each bridge transaction—a composite of a deposit (inflow) transaction on a source blockchain and a withdrawal (outflow) transaction on another—and b) identifying the value transferred in each such bridge transaction.
Figure 4: Total Inflow (ETH) - Total Outflow (SOL) Over Time: Wormhole Attack in Feb 2022.
3.1 Identifying Bridge Transactions

### v3

# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Token Bridges

@@ Abstract.
Between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen via attacks on “token bridges” — decentralized services designed to enable the transfer of a token between blockchains. While the individual exploits in each attack vary, a single design flaw underlies them all: the lack of end-to-end value accounting in cross-chain transactions. In this paper, we empirically analyze over 20 million transactions used by key token bridges during this period. We show that a simple invariant that balances cross-chain inflows and outflows is compatible with legitimate use, yet precisely identifies every known attack (and several likely attacks) in this data. Further, we show that this approach is not only sufficient for post-hoc audits, but can be implemented in-line in existing bridge designs to provide generic protection against a broad array of bridge vulnerabilities.
copyright: none
1. Introduction
@@
Careful accounting is key to the correct management of assets in virtually all financial systems. Indeed, since Pacioli popularized double-entry bookkeeping in the 14th century, it has been standard practice to separately track inflows (credits) and outflows (debits) to establish a net position — a balance sheet.
@@
Such accounting is implicit in blockchains as each and every transaction is recorded, immutable and in possession of implicit integrity. Any change in ownership for a given token is explicitly recorded via some past transaction and thus the net position for assets in a blockchain is generally consistent and well-known. However, these integral properties only hold within a single blockchain. As soon as one wishes to engage in transactions between chains (e.g., trading Ethereum on Solana), it requires building a system that steps outside these isolated environments and implements its own financial calculus between them, including its own accounting.
@@
Today, one of the principal mechanisms for such transactions is the cross-chain token bridge — a service using a pair of “smart contracts” (immutable programs stored on a blockchain) to synthesize inter-chain transactions that are not possible to express natively on a single chain. [^1] Some newer blockchain ecosystems, such as Cosmos and Avalanche, are themselves composed of multiple blockchains and offer some support for bridging across their own chains, but not with external chains. Such bridges are an extremely popular component of what is commonly referred to as “decentralized finance” (DeFi), and they routinely manage transaction volumes with value equivalent to \$US8–10 billion per month (DefiLlama, ).
@@
However, token bridges are just code. They can have bugs in their implementations, in the other services they depend upon, or in the mechanisms used to guard the integrity of their cryptographic secrets. As a result, attackers can, and do, exploit such vulnerabilities to extract significant value from such systems. For example, between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen in an array of attacks on token bridges.
The scale of these hacks has motivated a range of research focused on automating the detection of such vulnerabilities and attacks, through the analysis of low-level implementation details (e.g., using static analysis of smart contracts (Wang et al., 2024; Liao et al., 2024), predefined anomaly rules (Zhang et al., 2022), or the features of specific attacks (Lin et al., 2024)). However, virtually all of these efforts either require detailed modeling of each contract and bridge infrastructure (e.g., the complex protocol bridges run off-chain to relay messages between on-chain contracts), or are specialized to particular sets of vulnerabilities (e.g., fake deposit events (Lin et al., 2024)); they require continual updating, tuning and re-evaluation as they are tied to specific implementation artifacts.
In this work, we propose a complementary approach for detecting attacks on token bridges, by monitoring violations of the balance invariant — a measure of the difference between value inflows and value outflows at a token bridge. Unlike prior work, this approach is extremely simple to reason about because it focuses purely on detecting potential negative financial outcomes instead of focusing on the precise mechanism of attack that would lead to those outcomes. Indeed, the power of the approach arises from its technical simplicity. Because it abstracts away the complicated implementation details of bridge contracts (and the off-chain code interfacing with these contracts), the bridge invariant naturally captures attacks independent of the details of the vulnerabilities they target. We demonstrate this effectiveness by comprehensively surveying the twelve largest attacks (each \$US1m or more) between 2021 and 2023, and showing that while the details of their vulnerabilities vary considerably, all of them share the property that transactions are allowed to be “unbalanced” (i.e., that outflow can exceed inflow, less costs). We then empirically demonstrate, by auditing over 20m transactions from 11 different token bridges across 21 blockchains, that auditing transactions for this property is sufficient to retrospectively and automatically identify each of these past attacks.
We further show that such audits can be performed automatically and in real-time — either to alert bridge operators about potential fraud, or prevent such unbalanced transactions from completing, thus preventing any such loss.
@@
We argue that this approach is powerful both due to its simplicity (in a legitimate financial transaction, the value paid should be equivalent to the value received, less costs) and its independence from the vagaries of smart contract details. Moreover, it represents a straightforward mechanism that can prevent an entire class of existing attacks on bridges.
Our work directly inspired the design and implementation of the Bascule drawbridge security system, which has been used by Lombard—the largest Bitcoin liquid staking protocol—over the last year to secure the bridging of (over \$1B as of today) BTC deposited on Bitcoin to LBTC on Ethereum (Cubist, 2024).
@@
Figure 1. Cross-chain token bridging and the steps attackers can exploit to withdraw unbacked deposits.
2. Background
@@
We first give a brief introduction to smart contract blockchains and cross-chain token bridges. We then describe how token bridges, under attack, collapse in practice.
2.1. Smart Contract Blockchains
@@
Modern DeFi protocols (e.g., decentralized exchanges and lending protocols) are built on top of “smart contract” blockchains, like Ethereum. At their core, these chains, like the original Bitcoin blockchain, are distributed ledgers that manage accounts (public keys) and their balances (in native tokens like Ethereum’s ETH). Users interact with these chains by signing and broadcasting (to the distributed nodes that make up the chain network) transactions that, for example, transfer funds from their account (using the corresponding account private key) to another user’s account.
@@
Ethereum, and the smart contract chains it inspired since its release in 2015, differ from Bitcoin by extending the “simple” distributed ledger with a smart contract execution layer. A smart contract on Ethereum is an internal account—and, like a normal, externally owned account (EOA), it has a balance—that has associated code (EVM bytecode), which implements the smart contract’s program logic, and storage, which persists the program’s state across executions. Users interact with (i.e., execute) smart contracts much like they do when transferring funds from their account: they sign a transaction that encodes the smart contract to call (i.e., the contract address), the particular function to execute, and the arguments to call the function with. Instead of simply transferring funds from the user’s account, then, executing such a smart contract call transaction amounts to executing the smart contract bytecode—and any smart contracts the contract itself calls.
@@
Figure 2. Simplified USDC ERC-20 Token Contract with a Set of Common Functions.
2.2. ERC-20 Tokens
@@
One of the immediate applications of smart contracts—and to date still one of most popular—is to create custom tokens (or coins). To launch a new token, an organization no longer needs to launch a new blockchain; they can instead deploy a new contract that implements the ERC-20 token standard interface on a chain like Ethereum and take advantage of existing on-chain infrastructure like decentralized exchanges that make it easy to, for example, trade one kind of token for another—both native tokens (e.g., ETH) and other tokens. [^2] Most EVM chains, i.e., chains that use Ethereum’s execution layer, follow Ethereum standards like ERC-20. Most non-EVM chains like Solana (which have different execution models) have similar standards (e.g., SPL tokens in Solana’s case).
@@
Figure 2 shows a simplified variant of one such token contract—the USDC stablecoin contract. This contract tracks how many USDC tokens an account has and governs the spending of these tokens (much like a bank governs bank notes). For example, the contract’s mint function lets Circle (the company that owns the USDC contract) mint new tokens into a user’s account—e.g., after receiving the corresponding payment from the user off-chain (in US dollars). The contract exposes the ERC-20 interface that lets users (and smart contracts) transfer tokens from their account by simply calling functions like safeTransferFrom, and, in turn, use the tokens in any DeFi protocol (e.g., lending the USDC to different markets, exchanging USDC for ETH, etc.). Finally, the contract’s burn function “burns” a token out of circulation—and emits an event (Figure 3) that Circle’s off-chain code looks for before allowing a user to withdraw the corresponding USD fiat off-chain.
2.3. Cross-Chain Token Bridges
@@
While smart contracts make it easy to launch new tokens without spinning up new chains, there is no real shortage of new blockchains being deployed. Indeed, the modern blockchain ecosystem is a many-chain ecosystem. Blockchains like Avalanche, Base, and Solana have different design points—from cheaper “gas” execution costs, to higher throughput, lower latency, and different permission models—that make them better suited for different classes of application. This situation has resulted in applications that span many chains and cross-chain infrastructure that ultimately (try to) allow users to, for example, buy Dogwifhat NFTs on Solana using USDC tokens on Ethereum.
@@
Core to these applications and infrastructure—and the focus of our work—is the cross-chain token bridge. At a high level, a cross-chain token bridge makes it possible for a user to “transfer” their tokens from one chain (e.g., their USDC on Ethereum) to another chain (e.g., Solana) and then use the transferred tokens on the destination chain (e.g., on a Solana exchange trading USDC and Dogwifhat). [^3] While some cross-chain bridges do let users transfer one kind of token (e.g., USDC on Ethereum) for a completely different kind of token on the destination chain (e.g., Dogwifhat on Solana), these bridges essentially fuse the cross-chain token bridge with an exchange (or swap). We focus on token bridges not only because they are fundamental to other kinds of bridges, but also because they have higher volume, more liquidity, and more attacks. Since smart contracts cannot make network requests or otherwise access state outside their own storage, a cross-chain token bridge consists of smart contracts on both the source and destination chains, and off-chain infrastructure that serves to relay the “transfer” call across the two contracts.
Figure 3. Event emitted by an ERC-20 token (USDC).
@@
As Figure 1 shows, a typical cross-chain token transfer, i.e., a bridge transaction, consists of three phases:
@@
1. Deposit (on source chain). To initiate a cross-chain token transfer, the user first calls the bridge’s contract on the source chain:
@@
// ... validate the transfer request and the ERC-20 token contract
@@
This contract function—a simplified version of a bridge deposit function—processes the user’s deposit (step 2 in Figure 1) by validating the transfer request, e.g., against a list of supported tokens, and then transferring the user’s ERC-20 tokens to the bridge contract—recall that contracts are accounts with balances. Then, the function emits an event (step 3 in Figure 1) recording the user’s deposit details (including the deposit ID, the ERC-20 token contract address, the recipient on the destination chain, and value).
2. Off-chain relay. The emitted event is observed by an off-chain relayer in (step 4), which constantly monitors the source blockchain. The relayer first verifies the authenticity of the deposit event. If the event is authentic, the relayer then produces a signed receipt endorsing the deposit (step 5) and, typically, stores the receipt off-chain (step 6). Finally, this signed receipt is sent to the bridge’s withdrawal contract on the destination blockchain (step 7). Who submits the receipt varies across bridges—some bridges submit the receipt on the user’s behalf (in these cases, the signed receipt is simply a signed contract-call transaction), while others give users (and anyone willing to pay gas) the signed receipt and they, in turn, submit the receipt to the withdrawal contract to complete the transfer.
3. Withdraw (on destination chain). The withdrawal contract on the destination chain first processes the withdraw request by verifying the receipt then transfers the tokens to the recipient (step 8). In Chainswap’s withdraw case (simplified), for example, users call:
@@
This contract function first charges the caller a fee, then verifies the deposit details against receipt—both that the deposit was not already withdrawn and that the receipt signatures are valid—and finally transfers the tokens to the intended recipient. We consider a cross-chain transaction complete when the asset is released to the recipient (step 9).
We expect every cross-chain token bridge transaction to uphold the balance invariant: the value (and kind) of the tokens withdrawn—the outflow—should equal the value (and kind) of the tokens deposited—the inflow—minus the fees. In practice, they do not (Section 5).
3. Threat Model and Attack Vectors
In this section, we describe our threat model and the attack vectors in scope.
Threat Model. We consider attacks that seek to steal funds from a token bridge. The result of such attacks is that the token bridge is left in an unbalanced state, where the outflow of funds (withdrawals) exceeds the inflow of funds (deposits). In addition, we assume that the attacker always withdraws funds through the bridge’s designated withdraw function on the destination chain. Otherwise, the attacker can exploit any of the aforementioned components in Figure 1 or control the relayer (or submitter) keys. As we show later, our threat model captures the largest attacks on cross-chain bridges that have happened in the past.
Attack Vectors. We consider a broad array of attack vectors: attackers can exploit bugs in all three components—the deposit contract, the relayer, and the withdraw contract—and use stolen signing keys to steal funds. As we show below, these attack vectors effectively capture attacks that have happened in the past. Figure 1 highlights the precise steps in a cross-chain transaction that attackers have historically exploited, including:
@@ 3. Threat Model and Attack Vectors
    Bugs in the deposit contract. In step 2, the bridge contract verifies that it has received the correct amount of assets before emitting an event. Bugs in this verification logic could allow an attacker to deposit a smaller amount of assets than what is recorded by the bridge in the event (step 3). For example, Qubit’s deposit function did not properly validate the token address. This bug allowed an attacker to pass 0 for the token address, so the contract function did not actually transfer any funds from the attacker’s account but still emitted a Deposit event which allowed the attacker to withdraw actual tokens on the destination chain (rekt, ).
@@
    Bugs in the off-chain deposit verification. In step 5, the relayer verifies the authenticity of the deposit event emitted by the bridge contract, including whether the event is emitted by the bridge’s designated contract. Bridges that do not correctly verify deposits would allow attackers to withdraw assets that are never deposited.
@@
    Stolen relayer (or submitter) keys. In step 5, the relayer signs the deposit receipts which are then submitted to the withdrawal contract as evidence of a valid deposit. If the relayer key is compromised (e.g., as with the Ronin bridge (Thurman, 2022)) the attacker can forge a valid receipt and then withdraw assets that were never deposited by calling withdraw with the forged receipt.
    The same is true for bridges that submit receipts on behalf of users—and essentially restrict the withdraw callers to privileged submitter accounts. The bridge submitter keys (step 8) have similarly been compromised (e.g., as with AnySwap (CoinDesk, 2022)) and used to withdraw unbacked deposits.
@@
    Bugs in the withdraw verification. In step 8, the bridge contract verifies that the messages are signed by the relayer and have not been replayed. Bugs in this verification logic (e.g., as we saw with Wormhole (CertiK, 2022)) have allowed attackers to supply “valid” payloads that were not signed by the relayer and replay withdrawal requests with valid deposit receipts that have already been withdrawn.
We note that while our threat model captures a wide variety of vulnerabilities, it is not exhaustive. For example, we consider attacks that steal funds from a user (e.g., by tricking them into sending funds to an attacker-controlled address) out of scope. Similarly, we also exclude attacks that do not involve any component of a token bridge or do not have a withdrawal step (e.g., attacks that directly move funds from a token bridge’s wallet to an attacker’s wallet with a compromised key). Moreover, we do not consider swap bridges, which introduce additional complexity about the swap logic and the assets being swapped. Lastly, there can be attacks that withdraw fewer funds than deposited, and we consider these out of scope as bridges do not lose funds in such cases and are still in a balanced state.
4. Approach
The core hypothesis of our work is that value should be conserved within cross-chain transactions. That is, that the value of the asset inflow in such a transaction (i.e., the deposit) should equal the value of the asset outflow (i.e., the withdrawal). In token bridges, this invariant corresponds to a balancing of the inflow tokens and the outflow tokens (less any fees or transaction costs incurred by the bridge itself). When this balance invariant does not hold, it allows a range of opportunities for fraud, all of which involve greater outflows (withdrawals) than inflows (deposits).
Testing this invariant on a per transaction basis is straightforward in principle. However, since bridge transaction formats are not standardized, it requires a range of per-bridge and per-chain parsing in practice. Our methodology for normalizing this information focuses on two key pieces of information: a) identifying each bridge transaction—a composite of a deposit (inflow) transaction on a source blockchain and a withdrawal (outflow) transaction on another—and b) identifying the value transferred in each such bridge transaction.
4.1. Identifying and PairingToken Bridge Transactions